OpenStudy (anonymous):

Create two binomials first and then use the distribution method to find their simplified product. The simplified product is the answer.

hartnn (hartnn):

let the two binomials be x+a and x+b to find the product means to find (x+m)(y+n) now distributive property states (a+b)(c+d)=ab+ac+bc+bd now can u find (x+m)(x+n) ?

OpenStudy (anonymous):

Thank you, but they have to be legit numbers and I have no idea how to do this.

hartnn (hartnn):

if u want legitimate numbers then take m=n=1 but keep,x and y,they are required.

OpenStudy (anonymous):

I have no idea what your saying. =C

OpenStudy (anonymous):

The question is asking for you to make an equation with binomials and then solve the question you created

OpenStudy (anonymous):

this is an example The product of these two binomials is x2+x-20 Answer: (x+5)(x-4) but i cannot use this it must be unique

hartnn (hartnn):

so i made 2 binomials those are x+m and x+n

hartnn (hartnn):

ok,then lets take two binomials as (x+1)(x-1) ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now what?

hartnn (hartnn):

so using distributive property: (x+1)(x-1)=x^2+x-x-1 so what u get after simplification??

OpenStudy (anonymous):

um, x^2+ -1?

hartnn (hartnn):

almost correct :) its (x^2-1) after simplification.thats your final answer.

OpenStudy (anonymous):

Thank you so much! By any chance can we do just one more?

hartnn (hartnn):

welcome :) yes,by every chance.

OpenStudy (anonymous):

okay so, x^2+x-4 is this a binomial and how would we solve it

hartnn (hartnn):

no its not a binomial its a trinomial as it contains 3 different terms x^2,x,-4

OpenStudy (anonymous):

Oh, well what is another random binomial?

hartnn (hartnn):

have u solved quadratic equations earlier??

OpenStudy (anonymous):

I believe so, why?

hartnn (hartnn):

another random binomial would be 2x+7

OpenStudy (anonymous):

okay so 2x+7 how would I solve this

hartnn (hartnn):

because x^2+x-4 is a quadratic polynomial.

hartnn (hartnn):

u do not solve binomials or polynomials u solve binomial equations or polynomial equations 2x+7 is a binomial and solving it does not make sense 2x+7=0 is a binomial equation and it can be solved to get the value of x. so u want to solve this:2x+7=0 ??

OpenStudy (anonymous):

yes please

hartnn (hartnn):

so to isolate x first we subtract 7 from both sides to get 2x=-7 now since 2 is multiplied to x,we divide 2 on both sides x=-7/2 so u solved for x to get its value as -7/2 or -3.5 ok?

OpenStudy (anonymous):

ok so what is the answer

hartnn (hartnn):

to solve is to get the value of x so the final answer of our example is x=-3.5